### v1

Abstract
In recent years, machine learning models have seen widespread adoption across a broad range of sectors, including high-stakes domains such as healthcare, finance, and law. This growing reliance has raised increasing concerns regarding model interpretability and accountability, particularly as legal and regulatory frameworks place tighter constraints on the use of so-called black-box models in critical applications.
Although interpretable machine learning has attracted substantial attention, systematic evaluations of inherently interpretable models, especially for tabular data, remain relatively scarce and often focus primarily on aggregated performance outcomes.
To address this gap, we present a large-scale comparative evaluation of sixteen inherently interpretable methods, ranging from classical linear models and decision trees to more recent approaches such as Explainable Boosting Machines (EBMs), Symbolic Regression, and Generalized Optimal Sparse Decision Trees (GOSDT). Our study spans 216 real-world tabular datasets and goes beyond aggregate rankings by stratifying performance according to structural dataset characteristics, including dimensionality, sample size, linearity, and class imbalance. In addition, we assess training time and robustness under controlled distributional shifts.
Our results reveal clear performance hierarchies, particularly for regression tasks, where EBMs consistently achieve strong predictive accuracy. At the same time, we show that performance is highly context-dependent: Symbolic Regression and Interpretable Generalized Additive Neural Networks (IGANNs) perform particularly well in non-linear regimes, while GOSDT models exhibit pronounced sensitivity to class imbalance.
Overall, these findings provide practical guidance for practitioners seeking a balance between interpretability and predictive performance, and contribute to a deeper empirical understanding of interpretable modeling for tabular data.
keywords:
[…]
1 Introduction
The field of Interpretable Machine Learning (IML) has emerged, emphasizing models whose internal mechanisms are transparent and understandable to human users (molnar2020interpretable). This paradigm stands in contrast to the black box nature of complex architectures, such as deep neural networks, which prioritize predictive accuracy at the expense of explanatory insight (breiman2001statistical; rudin2019we).
It is also necessary to distinguish intrinsic interpretability from post-hoc explainability (often termed XAI). Post-hoc methods attempt to explain a complex model after it has been trained; however, because these are merely approximations, they may not accurately reflect how the model actually made a decision. In contrast, intrinsically interpretable models are transparent by design, guaranteeing that the explanation is an exact representation of the model’s internal logic (rudin_stop_2019). For this reason, such techniques are increasingly adopted in domains such as healthcare, finance, and law, where accountability and the understanding of the model decision-making process are essential (doshi2017towards; rudin_stop_2019). This demand is further amplified by evolving regulatory frameworks, such as the European Union’s AI Act and the General Data Protection Regulation (GDPR), which mandate stringent transparency requirements and the right to explanation for high-risk automated systems. In many tabular-data applications, intrinsically interpretable models, such as linear models, sparse decision trees, or symbolic models, are already regarded as strong candidates due to their clarity and ease of validation by domain experts (zeng2017interpretable; la2021contemporary). Despite this growing practical importance, empirical guidance on the comparative performance of interpretable models remains limited.
Most existing benchmarking studies focus on general-purpose machine learning algorithms rather than interpretable ones (fernandez2014we; wainer2016comparison; olson2017pmlb; scholz_comparison_2021). As a result, practitioners often lack systematic evidence about (i) whether some interpretable models tend to perform better than others across diverse real-world datasets, and (ii) how model performance changes as a function of dataset characteristics such as dimensionality, sample size, linearity, or class imbalance. A few recent works examine specific subclasses of interpretable models, such as Generalized Additive Models (GAMs) (kruschel2025challenging), but do not cover the broader landscape of modern IML approaches nor analyze performance variation under different data conditions. Moreover, comparative evaluations are typically conducted on small collections of hand-selected datasets or on synthetic data, which may limit generalizability.
At the same time, the ecosystem of interpretable methods has expanded substantially over the past decade. Techniques such as Explainable Boosting Machines (EBMs) (nori2019interpretml), Interpretable Generalized Additive Neural Networks (IGANNs) (kraus2023interpretable), Generalized Optimal Sparse Decision Trees (GOSDTs) (mctavish2022fast), and modern Symbolic Regression methods (la2021contemporary) provide new modeling options whose empirical properties have not yet been systematically compared. Consequently, users are faced with an increasingly fragmented methodological landscape without clear, data-driven recommendations for model selection.
Our contribution
To address these gaps, we present what is, to the best of our knowledge, the most extensive comparative study of intrinsically interpretable machine learning models to date [^1] The supplementary material and the code required to reproduce all experiments are available at: https://github.com/mattiabilla/CA-IML.
  • Using the full Penn Machine Learning Benchmark (PMLB) repository (olson2017pmlb), we evaluate 16 interpretable models across 137 regression and 79 classification datasets. Our analysis includes both traditional and recent IML methods with stable, well-maintained implementations (9 regression models and 7 classification models), assessing both predictive performance and training time. Our experimental design includes 4-fold cross-validation, systematic hyperparameter optimization, and rigorous preprocessing, resulting in over 14,000 model evaluations and more than 700,000 individual training runs.
  • Beyond aggregated results, we investigate how model performance varies with key dataset characteristics, including structural properties (e.g., number of features and samples), degree of linearity, and class imbalance (maciel_measuring_2016; lorena_how_2020). This stratified analysis provides deeper insights into the conditions under which different interpretable models excel or underperform, offering practical guidance for both researchers and practitioners.
  • We explicitly evaluate generalization robustness by contrasting standard in-sample performance with out-of-sample settings. This approach simulates the phenomenon of distributional shift (covariate or label shift), allowing us to assess how well models maintain their performance when the testing distribution diverges from the training data.
Our empirical analysis reveals several important patterns. For regression tasks, clear performance hierarchies emerge under different data regimes: methods such as EBM and Symbolic Regression (SR) consistently excel on non-linear or low-sample problems, whereas regularized models like LASSO and Generalized Linear Models (GLMs) demonstrate superior robustness against distributional shifts and high dimensionality compared to standard Linear Regression (LR). In contrast, for classification, performance differences between models are less sharply stratified: most methods behave similarly across data conditions, making it more difficult to identify a preferable model a priori. This weaker stratification is partly attributable to the smaller number of available classification datasets in PMLB, but it also suggests that classification performance may be inherently less sensitive to the dataset metrics considered here.
The remainder of this paper is organized as follows. Section 2 reviews related work on benchmarking with a focus on interpretable models. Section 3 details the experimental setup, including the selection of machine learning models and dataset preprocessing. Section 4 specifies the evaluation protocol, defining the cross-validation design, generalization scenarios, and performance metrics. Section 5 describes the dataset metrics used for stratification and the statistical procedures for model ranking. Sections 6 and 7 present and analyze the empirical results for the regression and classification tasks. Section 8 relates our results to existing work and discusses their implications. Section 9 concludes the paper and highlights directions for future research.
2 Related Work
A number of studies in the literature have conducted comparative analyses of machine learning models (fernandez2014we; wainer2016comparison; olson2017pmlb; zhang2017up; scholz_comparison_2021). However, the majority of these works do not focus specifically on IML models. While interpretable models are sometimes included in broader comparisons, recent state-of-the-art approaches such as SR (la2021contemporary), EBM (nori2019interpretml), or GOSDT (hu2019optimal; mctavish2022fast) are often omitted.
A recent IML-focused work, kruschel2025challenging, presents a systematic comparison of interpretable models with a focus on GAMs. Their analysis provides strong evidence against the presumed trade-off between accuracy and interpretability. Nevertheless, the scope of their study remains restricted, as other relevant interpretable approaches are not considered.
Moreover, most existing comparative evaluations, such as the work by fernandez2014we, rely on aggregated performance results across a collection of datasets, rather than systematically studying how model performance varies under different data conditions. As a result, important nuances in model behavior may be obscured. In contrast, scholz_comparison_2021 explicitly investigate how classification methods respond to varying structural dataset characteristics, such as dimensionality and target distributions, using synthetic data to evaluate performance under controlled conditions. Their investigation reveals that small heterogeneous ensembles, specifically simple averages of just 3 to 5 models, generally yield the best predictive performance. They further identify that specific settings favor certain algorithms: Bagged CART excels in low-dimensional, high-sample scenarios, while nearest shrunken neighbor classifiers are most effective for unbalanced datasets. This type of analysis underscores the importance of accounting for dataset properties in comparative studies, a perspective that remains largely unexplored in the context of IML. While synthetic datasets enable controlled experimentation, they may fail to fully capture the complexity, heterogeneity, and noise patterns found in real-world data, which can limit the external validity of the conclusions regarding which specific models are optimal for a given complexity scenario
A related approach using real-world datasets was presented by kraus2023interpretable, who stratified datasets from the PMLB repository by sample size and dimensionality. However, other critical aspects, such as dataset linearity, and the complexity of decision boundaries, remain insufficiently explored.
Another recurring limitation in many comparative evaluations is the restricted choice of benchmark datasets, often confined to a small number of manually selected examples, typically drawn from well-known repositories such as UCI (uci2019). This raises concerns regarding the generalizability and robustness of the reported findings. Moreover, rigorous statistical testing is frequently absent, further limiting the reliability of conclusions. As argued by herrmann2024position, empirical machine learning research is often presented as confirmatory, even though comparative evaluation should instead be treated as exploratory. Comparative analyses are also commonly conducted in the context of introducing new models, which can unintentionally bias results in favor of the proposed method.
In contrast to these practices, and to prior studies that either relied on narrow subsets of datasets (kruschel2025challenging) or on synthetic data designed to isolate specific conditions (scholz_comparison_2021), our evaluation will draw on a broad set of real-world datasets from PMLB, excluding only those unsuitable for our setting (see Section 3.2). This approach mitigates the risk of selection bias and ensures a more comprehensive assessment of model behavior across diverse problem types. Furthermore, to move beyond aggregate performance and gain insight into how models behave under different data conditions, we stratify results using several dataset-level metrics (see Section 5.1), enabling a more fine-grained and interpretable analysis and test generalization properties both in- and out-of-sample.
3 Experimental Setup
This section describes the experimental setup underlying our large-scale benchmark of IML models. We first present the set of candidate models included in the study and motivate their selection based on interpretability and implementation maturity. We then detail the dataset collection process and preprocessing pipeline applied to construct a consistent benchmark. Finally, we describe the computational environment and the hyperparameter search spaces and budgets used to ensure a fair and reproducible comparison across methods.
3.1 Selection of ML Models
Method Acronym Task Reference Library
Linear Regression LR REGR scikit-learn scikit-learn1
Logistic Regression LR CLF scikit-learn scikit-learn
Lasso Regression LASSO REGR scikit-learn scikit-learn
Generalized Linear Models GLM REGR scikit-learn scikit-learn
Polynomial Regression w/ LASSO PR+LASSO REGR scikit-learn scikit-learn
Explainable Boosting Machine EBM CLF, REGR nori2019interpretml InterpretML2
Interpretable Generalized Additive Neural Network IGANN CLF, REGR kraus2023interpretable IGANN3
Decision Trees DT CLF, REGR scikit-learn scikit-learn
Generalized Optimal Sparse Decision Trees GOSDT CLF mctavish2022fast GOSDT4
Naive Bayes NB CLF scikit-learn scikit-learn
k-Nearest Neighbors k-NN CLF, REGR scikit-learn scikit-learn
Symbolic Regression SR REGR cranmer2023interpretable PySR5
Table 1: Overview of the interpretable models included in our comparative evaluation. For each method, we report its acronym, the type of task it addresses (regression and/or classification), key references, and the software library or implementation used in our experiments.
For our comparative evaluation, we prioritize IML models with stable and publicly available implementations. This criterion is crucial to ensure that our results are both reproducible and relevant to practitioners deploying these models in real-world scenarios. A complete summary of the methods included in this study, along with their respective implementations, is provided in Table 1.
In total, we examine nine regression and seven classification methods. We selected these methods to cover a broad spectrum of interpretability and expressiveness, spanning various modeling strategies (e.g., linear, probabilistic, instance-based). Our selection includes both established, widely investigated baselines (e.g., Linear and Logistic Regression, Decision Trees, k-NN) and novel approaches (e.g., Symbolic Regression, Optimal Sparse Decision Trees).
The baseline of our analysis is formed by the family of linear models (Linear/Logistic Regression, Lasso, and Polynomial Regression). While these methods offer the highest degree of transparency, their expressiveness is limited by their fixed functional form. Increasing the level of expresiveness, we also consider Generalized Additive Models. Within this class, we specifically selected Explainable Boosting Machines (nori2019interpretml) and Interpretable Generalized Additive Neural Networks (kraus2023interpretable). We focused on these two methods because recent comparative work by kruschel2025challenging positions them on the performance–interpretability Pareto frontier among modern GAM variants; furthermore, unlike many experimental alternatives, they offer mature and well-documented implementations.
To explore a further level of flexibility beyond the additive constraints of GAMs, we include Symbolic Regression. Unlike standard parametric or semi-parametric methods, SR does not assume a pre-defined model structure; instead, it searches the space of mathematical expressions to discover the underlying functional form of the data. Given that SR is computationally demanding, we limit our evaluation to a single representative: PySR (cranmer2023interpretable). We selected PySR over alternative implementations (such as Operon (operon) or QLattice (brolos2021approach)) as recent benchmarks demonstrate that it delivers highly competitive accuracy–interpretability trade-offs (de2024srbench++), alongside a mature and efficient Python implementation.
Finally, regarding tree-based methods, we include both classical greedy Decision Trees (DTs) and Generalized Optimal Sparse Decision Trees (mctavish2022fast). We include GOSDT because it employs a bound-driven search for provably optimal trees on binary variables. This approach results in smaller, more interpretable trees compared to greedy splitting methods used in standard DTs, while maintaining competitive predictive performance.
A detailed description and discussion of these methods is provided in Appendix A.
3.2 Dataset Selection and Preprocessing
We base our study on the Penn Machine Learning Benchmark (PMLB) repository (olson2017pmlb), a curated collection of 450 datasets that has become a widely used benchmark suite, and is frequently used in studies on interpretable models, especially on SR (la2021contemporary; de2024srbench++). PMLB contains 179 classification datasets and 271 regression datasets. Not all of these are suitable for our purposes, so we apply a sequence of filtering and preprocessing steps summarized in Figure 1, where C and R denote the number of classification and regression datasets retained after each step. Orange boxes indicate filtering (dataset removal), while yellow boxes denote transformations applied to the datasets.
We first remove datasets tagged as “deprecated” as well as synthetic datasets. Many synthetic PMLB datasets are derived from exact physics-based equations, or similarly specialized mechanisms; including them may bias performance rankings toward models that can closely adapt to these specific functional forms, and are less representative of typical real-world applications where the generative process is unknown.
For classification, we restrict attention to binary problems. This choice reflects structural constraints of some of the interpretable models considered and simplifies the construction of out-of-sample scenarios described in Section 4.2. Conceptually, this does not limit generality, as multiclass problems can be decomposed into multiple binary tasks.
Since most interpretable methods do not natively handle missing values, and imputation can distort feature distributions in a method-dependent way, we rely on simple deletion. First, we drop columns with more than 30% missing values, retaining only those datasets where at least half of the original columns remain. Next, we remove rows containing any missing values, keeping only datasets that preserve at least 50% of the original rows. Datasets requiring more extensive deletion are discarded, as they would be too heavily altered and could bias the results.
After these selection steps, 216 datasets remain: 137 regression datasets and 79 binary classification datasets.
For each dataset, categorical variables are transformed via one-hot encoding and numerical features are scaled using median/IQR scaling, which reduces sensitivity to outliers (RobustScaler from scikit-learn). The scaler is fitted only on the training split of each dataset, and the learned parameters are then applied to the corresponding validation and test sets to avoid information leakage. Figure 2 reports the distribution of the number of samples and features for the final set of selected and preprocessed datasets.
Figure 1: Dataset filtering and missing values handling. Orange boxes indicate filtering steps where datasets are removed based on minimum quality criteria, while yellow boxes represent transformation steps where datasets are modified but retained.
Figure 2: Number of samples and features of the selected datasets.
3.3 Hyperparameter Optimization
Hyperparameter optimization is necessary to ensure that performance comparisons are not confounded by arbitrary configuration choices. We therefore tune hyperparameters independently for each model–dataset combination.
When available, the search space was initialized based on the hyperparameter ranges suggested in the papers introducing the corresponding methods. Hyperparameter ranges were further constrained to promote smaller and more comprehensible models, in line with interpretability objectives. For example, for Decision Trees, the maximum tree depth was limited to 6, as deeper trees would typically result in hundreds or thousands of nodes, making them difficult to interpret meaningfully. The full list of hyperparameters and their corresponding search ranges used in our experiments is reported in Table 3 in Appendix B. The hyperparameter optimization objective was the maximization of the performance measures described in Section 4.3.
Hyperparameter search was conducted using the Tree-structured Parzen Estimator (TPE) algorithm (bergstra2011algorithms), as implemented in the Optuna Python library (akiba2019optuna). This Bayesian optimization method allows for a more efficient exploration of the hyperparameter space compared to random search, which is particularly advantageous for high-dimensional configurations in some of the tested models. We executed at most 50 optimization trials per setting, limiting the total optimization time to 6000 seconds, corresponding to an average of approximately 2 minutes per trial. For models with higher computational demands, such as SR and GOSDT, this time constraint was explicitly enforced to ensure a feasible evaluation.
In this study, which involves more than two hundred datasets, we balance the need for reliable hyperparameter selection with the computational constraints imposed by large-scale experimentation. To this end, we adopt a K-fold cross-validation strategy for each model–dataset combination. Hyperparameters are optimized on the validation subset of each fold using a dedicated search procedure, rather than via fully nested cross-validation, in order to keep the computational cost tractable. While less expensive than a nested scheme, this approach still provides a robust and unbiased estimate of model generalization performance.
After selecting the optimal configuration for a given fold, models are retrained on the combined training and validation data under a fixed time budget of 300 seconds, and evaluated on the held-out test set. Final performance scores are obtained by aggregating test-set results across the four folds using the median.
3.4 Compute environment
All experiments were executed on a shared compute cluster managed by SLURM. Each node was equipped with two Intel Xeon Gold 5320 processors (2.2 GHz), providing a total of 52 cores and 512 GB of shared memory. To guarantee a fair comparison, every individual job was allocated a single CPU core and 16 GiB of RAM, without GPU acceleration. This strict standardization enhances reproducibility and prevents discrepancies arising from differing multiprocessing or threading implementations that could otherwise bias the results.
4 Evaluation protocol
This section specifies the evaluation protocol used to assess model performance across datasets. We define a cross-validation design and an experimental framework to probe model robustness in both standard settings and in the presence of controlled distributional shifts. To provide a rigorous foundation for these comparisons across diverse tasks, we first establish a common notation for our data structures and predictive objectives.
Formally, we denote each dataset as $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, where $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents the $d$-dimensional vector of processed features for the $i$-th sample, and $y_{i}$ denotes the corresponding ground truth target. This framework encompasses both classification and regression; the target domain is defined as $y_{i}\in\{0,1\}$ for the classification task and $y_{i}\in\mathbb{R}$ for regression task. We let $\hat{y}_{i}$ denote the prediction produced by the evaluated methods for the input $\mathbf{x}_{i}$. Using this notation, the following subsections detail how these data are partitioned and manipulated to simulate various testing environments.
4.1 Cross-validation design
We adopt a 4-fold cross-validation scheme to balance robustness of performance estimation with computational efficiency. This provides sufficient variability across data partitions to explore different training and validation configurations, while remaining sustainable for small datasets and large-scale benchmarking.
Within each fold, data are split into training (50%), validation (25%), and test (25%) subsets. This protocol is used in the in-sample experiments both for hyperparameter optimization on the validation splits and for performance estimation on the corresponding test splits. For out-of-sample experiments, hyperparameters are similarly optimized on the validation split and are not re-tuned, ensuring that performance under distributional shift is assessed without additional adaptation.
4.2 In-sample and out-of-sample generalization
In our evaluation, we distinguish between two critical experimental scenarios: in-sample and out-of-sample settings. The in-sample setting represents the standard evaluation scenario where the training and test sets are drawn from the same underlying data distribution (e.g., a simple random split), allowing us to assess a model’s performance on data statistically identical to what it was trained on. In contrast, the out-of-sample setting simulates the more challenging real-world phenomenon of distributional shift (also known as covariate shift or label shift), where the training and test sets exhibit different statistical properties.
To systematically assess the in- and out-of-sample generalization capabilities of the evaluated models across both regression and binary classification tasks, we designed a two-stage evaluation protocol ensuring statistical consistency within samples and controlled distributional shifts across samples.
In-sample Generalization
In-sample generalization was evaluated using stratified sampling to preserve the distributional properties of the target variable $y$ across the training and test subsets.
  • Classification. The stratification was performed with respect to the binary target variable, ensuring that class proportions remained consistent across splits.
  • Regression. Since the target is continuous, we partitioned $y$ into deciles and performed stratified sampling over these bins to ensure that both subsets reflected the overall distribution of target values.
Out-Of-Sample Generalization
To evaluate model robustness under distributional shifts, we constructed test sets with deliberately altered target distributions.
  • Classification. To introduce class-imbalance shifts, we modified the fraction of positive samples in the test set. Let $f$ denote the test fraction and $\alpha_{\mathcal{D}}$ the overall positive rate. We constrain the test proportion $\alpha$ to lie in the interval
    $\frac{l}{f}\alpha_{\mathcal{D}}\leq\alpha\leq\frac{u}{f}\alpha_{\mathcal{D}},$
    where $l$ and $u$ denote lower and upper bounds on the share of positives assignable to the test set. In our experiments, we set $f=1/4$, $l=1/8$, and $u=1/2$, yielding shifts of up to twofold differences in class proportions between training and testing subsets. To ensure strong distributional change between training and test sets, we require that
[…]
    so that the balancing between training and test set gets reverted, when possible.
  • Regression. We simulated out-of-distribution conditions by systematically excluding one quartile of the target distribution to form the test set. Each quartile was omitted once in a fourfold procedure, ensuring that each portion of the distribution was tested independently.
This procedure produces controlled target shifts that enable a fair assessment of each model’s generalization under distributional changes, while preserving sufficient statistical diversity within each dataset.
4.3 Performance Metrics
To evaluate model performance, we primarily focus on the coefficient of determination ($R^{2}$) for regression and the F1 score for classification. These metrics were selected as the most representative measures of predictive quality for our analysis. For regression, $R^{2}$ provides a normalized measure of goodness of fit that accounts for the variance (and implicitly, the noise) in the target variable. It is defined as:
$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}},$
where $\bar{y}$ denotes the mean of the observed targets.
For classification, the F1 score offers a balanced assessment of precision and recall, making it particularly suitable for datasets where class distribution may be imbalanced. The F1 score is computed as:
$\text{F1}=2\cdot\frac{\text{precision}\cdot\text{recall}}{\text{precision}+\text{recall}}.$
We also computed several other standard metrics, such as Mean Squared Error (MSE) and Mean Absolute Error (MAE) for regression, as well as precision, recall, and accuracy for classification. Because these yielded results highly correlated with our primary metrics, we focus our discussion on $R^{2}$ and F1 to maintain clarity and avoid redundancy. Detailed results for all other metrics are provided in the supplementary material.
5 Model Evaluation
5.1 Dataset Metrics for Stratification
To better understand how IML models behave across heterogeneous problems, we characterize each dataset $\mathcal{D}$ using a small set of complexity descriptors that were proposed for classification (lorena_how_2020) and regression (maciel_measuring_2016). These descriptors summarize size, sparsity, class/target structure, and local neighborhood overlap, and are subsequently used to stratify performance across the PMLB benchmark.
For classification problems, let $\mathcal{C}$ denote the set of classes and $p_{c}$ the empirical proportion of class $c\in\mathcal{C}$. We compute metrics belonging to the following classes.
Size and sparsity
We characterize dataset size and sparsity using the following indicators:
  • Number of samples ($N$).
  • Number of features ($d$).
  • Average number of features per point (T2): $\text{T2}=d/N$.
  • Average number of PCA dimensions per point (T3):
    $\text{T3}=d^{\prime}/N\,$,
    where $d^{\prime}$ is the number of principal components explaining $95\%$ of the variance in $\mathbf{X}$.
These indicators jointly describe the scale and effective sparsity of a dataset. Larger values of $N$ typically stabilize empirical performance estimates, whereas small-sample regimes increase sensitivity to noise and model specification. High dimensionality ($d\gg 1$) can favor models with built-in regularization or feature selection while penalizing methods relying on local neighborhoods. The ratios T2 and T3 explicitly capture sparsity by relating dimensionality to sample size: large values indicate that data occupy a high-dimensional space relative to the number of observations, a regime in which decision boundaries are poorly supported and both parameter estimation and generalization become more challenging. In particular, T3 approximates intrinsic dimensionality after removing linear correlations, highlighting cases where many effective degrees of freedom must be learned from limited data.
Classification-specific structure
We characterize class separability, local overlap, and class balance using the following indicators:
  • Maximum Fisher’s discriminant ratio (F1):
    $\text{F1}=\frac{1}{1+\max_{j}r_{j}}\in[0,1]\,;$
    where
    $r_{j}=\frac{\sum_{c\in\mathcal{C}}p_{c}(\mu_{cj}-\mu_{j})^{2}}{\sum_{c\in\mathcal{C}}p_{c}\sigma_{cj}^{2}}\,$
    and $\mu_{cj}$ and $\sigma_{cj}^{2}$ are the class $c\in\mathcal{C}$ conditional mean and variance of feature $j=1,\dots,d$, and $\mu_{j}$ is the global mean.
  • Error rate of the 1-nearest neighbor classifier (N3):
    $\text{N3}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\bigl(y_{i}^{\text{NN}}\neq y_{i}\bigr)\in[0,1].$
    where $y_{i}^{\text{NN}}$ is the label of the nearest neighbor of sample $\mathbf{x}_{i}$.
  • Entropy of the 5 nearest neighbors (E5):
    $\text{E5}=-\frac{1}{N}\sum_{i=1}^{N}\left[p_{i}\log p_{i}+(1-p_{i})\log(1-p_{i})\right]\in[0,1]\,$
    where
    $p_{i}=\frac{1}{5}\sum_{j\in\mathcal{N}_{5}(i)}\mathbb{I}(y_{j}\neq y_{i})\,$,
    and $\mathcal{N}_{5}(i)$ is the set of the five nearest neighbors of $\mathbf{x}_{i}$.
  • Entropy of class proportions $C1_{\text{c}}$:
    $\text{C1}_{\text{c}}=-\frac{1}{\log C}\sum_{c\in\mathcal{C}}p_{c}\log p_{c}\in[0,1].$
    where $p_{c}$ is the proportion of class $c$ in the dataset and $C=|\mathcal{C}|$ the number of classes.
These indicators jointly capture different aspects of classification difficulty. The Fisher-based measure $F1$ quantifies the extent to which at least one feature can separate classes through a simple threshold. The numerator of $r_{j}$ captures between-class variability, while the denominator captures within-class variability. Thus, low F1 values indicate near-axis-aligned separability and higher values reflect increasingly entangled class structure. The neighborhood-based metrics N3 and E5 probe local class overlap at different spatial scales: high values indicate irregular or ambiguous decision boundaries where nearby samples frequently belong to different classes. Finally, $\text{C1}_{\text{c}}$ measures global class balance, highlighting regimes in which skewed label distributions may disproportionately affect certain learning algorithms. Together, these descriptors allow us to stratify performance by separability, locality, and imbalance.
Regression-specific structure
We characterize functional simplicity, linearity, and local smoothness using the following indicators:
  • Maximum feature–target correlation $C1_{\text{r}}$
    $\text{C1}_{\text{r}}=\max_{j=1,\dots,d}|\rho(x_{j},y)|\in[0,1].$
    where $\rho(x_{j},y)$ denotes the Spearman correlation between feature $j$ and the target.
  • Distance to a linear model (L1):
    $\text{L1}=\frac{1}{N}\sum_{i=1}^{N}|y_{i}-f(\mathbf{x}_{i})|,$
    where $f$ is a fitted multiple linear regressor on the dataset $\mathcal{D}$.
  • Error of a 1-nearest neighbor regressor (S3):
    $\text{S3}=\frac{1}{N}\sum_{i=1}^{N}\bigl(\widehat{y}_{i}^{\text{NN}}-y_{i}\bigr)^{2},$
    where $\widehat{y}_{i}^{\text{NN}}$ denotes the leave-one-out 1-NN prediction.
These metrics capture complementary aspects of regression difficulty. The correlation-based measure $\text{C1}_{\text{r}}$ quantifies whether the target is strongly driven by a single feature, indicating low intrinsic complexity and near one-dimensional structure. The linearity measure L1 assesses how well the data conform to a global linear model, with low values signaling simple relationships or low noise levels, and higher values reflecting the need for increased model expressiveness. Finally, the neighborhood-based error S3 probes local smoothness of the regression surface: high values indicate that nearby points may have substantially different outputs, suggesting irregular functions or sparse sampling of the input space.
Figure 3: Correlation plot of the metrics for classification datasets computed using the Spearman correlation.
Figure 4: Correlation plot of the metrics for regression datasets computed using the Spearman correlation.
The set of metrics considered, summarized in Table 2, spans multiple dimensions of dataset complexity. Their complementarity is assessed by examining pairwise Spearman correlations, reported in Figures 3 and 4. The predominantly low correlations across most pairs suggest these descriptors provide complementary information, justifying their joint use in our stratification. Notable exceptions include the expected structural dependencies between T2 and T3, as well as their natural correlation with $N$ and $d$. In classification, we observe a moderate correlation between N3 and E5, as both metrics probe local neighborhood purity. Similarly, for regression, $\text{C1}_{\text{r}}$ and L1 show a distinct correlation, reflecting their shared sensitivity to the linearity of the feature-target relationship.
In the following, we use these metrics to partition the PMLB benchmark into strata and analyze whether the relative performance of interpretable methods remains stable across these regions or changes substantially under more complex data conditions.
Metric Measured Property Complexity ($\nearrow$) Tasks
NN Dataset size $\searrow$ Class., Reg.
$d$ Dimensionality $\nearrow$ Class., Reg.
T2 Data sparsity $\nearrow$ Class., Reg.
T3 Intrinsic data sparsity $\nearrow$ Class., Reg.
F1 Single feature discrimination $\nearrow$ Class.
N3 Local class overlap $\nearrow$ Class.
E5 Local class overlap $\nearrow$ Class.
$\text{C1}_{\text{c}}$ Class Balance $\searrow$ Class.
$\text{C1}_{\text{r}}$ Single feature discrimination $\searrow$ Reg.
L1 Linearity level $\nearrow$ Reg.
S3 Smoothness / Sampling quality $\nearrow$ Reg.
Table 2: Complexity Metrics for Classification and Regression.
5.2 Model Ranking
For each dataset $\mathcal{D}$, models were ranked according to their predictive performance, with a rank of 1 assigned to the method exhibiting the best performance. We then calculated the average rank for each model either over the entire dataset collection or stratified by specific evaluation metrics introduced in Section 5.1. To formally compare the methods and assess the statistical significance of the observed differences in their average ranks, we employed the Friedman test (friedman1937use). This non-parametric test is appropriate for comparing multiple models across multiple datasets because it operates on ranks rather than raw performance scores. This is crucial as performance metrics across diverse datasets often violate the normality assumption and homoscedasticity required by parametric tests like ANOVA. If the null hypothesis (that all models perform equally well) was rejected by the Friedman test, we proceeded with the Nemenyi post-hoc procedure (nemenyi1963distribution) using a significance level of $\alpha=0.05$. The Nemenyi test determines the critical difference required for two average ranks to be considered statistically distinct, allowing us to identify groups of methods that are statistically indistinguishable from one another in terms of performance.
6 Regression Results
In this section, we present the results of our experiments for the regression task. Our analysis begins with an evaluation of the methods’ relative performance aggregated across the entire dataset collection in the in-sample setting. The results are then analyzed in greater detail by stratifying them according to the dataset metrics introduced in Section 5.1. Finally, we compare the methods’ performance in the Out-of-Sample configuration and examine their computational training time.
6.1 Regression Performance (In-Sample)
Figure 5: Ranking comparison for the regression tasks between in-sample and out-of-sample settings.
Figure 5 shows the relative ranking of the regression methods based on the $R^{2}$ metric. We introduce a modified version of the critical-difference diagram in which brackets denote groups of models whose rankings are not statistically distinguishable, indicating comparable predictive performance. Models are ordered from bottom to top according to their average rank, from better to worse.
Figure 6: Comparison of model complexity, measured by the number of non-zero linear terms, between Linear Regression, LASSO, and Polynomial Regression with LASSO.
For the in-sample setting (left-hand side), we observe three clearly defined groups of methods. The EBM model outperforms the others by a substantial margin. The second group includes Polynomial Regression with LASSO regularization, IGANN, and Symbolic Regression. Notably, Symbolic Regression, typically valued for producing highly interpretable mathematical expressions, achieves performance comparable to that of GAMs, which are typically more complex to interpret. The final group comprises the remaining models, primarily linear regressors, Decision Trees, and k-NN, which generally exhibit lower performance.
Quantifying interpretability is generally architecture-specific, precluding a direct comparison across heterogeneous models (e.g., k-NN versus Symbolic Regression). However, within the family of linear models, we can structurally compare interpretability by analyzing the sparsity of the resulting equations. To quantify this sparsity, we apply a numerical threshold, treating any coefficient with an absolute value below $10^{-6}$ as zero. Figure 6 illustrates the distribution of the remaining non-zero coefficients, showing that LASSO effectively minimizes model complexity compared to standard Linear Regression. In contrast, PR+LASSO exhibits a marked increase in the number of terms due to the expansion of the feature space. Beyond the sheer number of terms, the inclusion of higher-order transformations and interaction effects makes Polynomial Regression inherently less transparent than purely linear models. At the same time, this added structural complexity enables the model to represent non-linear relationships that are fundamentally inaccessible to linear formulations. Therefore, PR+LASSO exemplifies a clear performance–interpretability trade-off: increased expressive power can yield improved predictive accuracy, but results in symbolic representations that are more elaborate and require greater effort to interpret.
6.1.1 Stratified Results for Regression
Figure 7: Stratified results for the regression task (in-sample).
In the previous sections, we presented the aggregated results across all 137 regression datasets. We now extend the analysis by stratifying the results according to the dataset metrics introduced in Section 5.1, allowing us to capture more nuanced behaviors of the regressors.
Figure 7 presents the stratified results, where each plot corresponds to one metric and divides the datasets into three bins representing the terciles. When multiple datasets share a metric value that lies at the boundary between two bins, those datasets are included in both bins.
When considering the number of samples, we observe an improvement in the performance of k-NN, as expected, and of Decision Trees as the number of samples increases. Conversely, Symbolic Regression exhibits a deterioration in relative performance. In this regime, the selected SR library (PySR) enables batching to make the symbolic model search more efficient, as suggested in the official documentation [^2] https://astroautomata.com/PySR/v1.5.9/api. Despite this, Symbolic Regression appears to struggle in such settings, possibly due to inefficiencies in the symbolic search process when handling large and potentially noisy datasets. This behavior is also reflected in the T2 and T3 metrics, where the relative performance of SR decreases at lower values of these metrics.
The opposite trend is observed when considering the number of features. Surprisingly, SR seems relatively insensitive to higher dimensionality, contrary to common expectations. This may be explained by the feature selection step performed by PySR, which relies on a random forest to reduce dimensionality before model search.
As expected, in high-dimensional spaces, instance-based methods such as k-NN suffer from the curse of dimensionality. PR+LASSO is also affected, though to a lesser extent, as the polynomial expansion increases the number of features, thereby magnifying the impact of high dimensionality.
When analyzing the linearity of the problem using L1, we find that when datasets are well described by a linear model (i.e., low L1 values), nearly all regressors, except for DTs and k-NN, achieve comparable performance. As non-linearity increases, methods capable of modeling non-linear relationships, such as SR, EBM, and IGANN, outperform the others. PR+LASSO, despite being a linear combination of polynomial terms, appears less sensitive to increasing non-linearity, likely due to its flexible basis expansion.
For the local noise metric, S3, derived from the error of a 1-NN regressor, we do not observe strong or unexpected patterns, except in the highest noise regime, where k-NN surprisingly does not perform the worst. This may indicate that other regressors are similarly affected by noise, reducing relative differences.
In datasets containing a feature that is highly correlated with the target (higher values for metric $\text{C1}_{\text{r}}$), all regressors perform more similarly. When the features are weakly correlated, we observe the expected distribution of model performances, with k-NN and Decision Trees outperforming linear models.
Finally, we note that the various linear regression variants perform comparably overall, except for LASSO, which shows advantages in high-dimensional settings (metrics T2, T3, and number of columns), confirming its intended regularization behavior.
6.2 Out-Of-Sample Regression
The previous sections presented the results obtained in the in-sample setting. To further deepen our analysis, we now examine the results derived from the out-of-sample simulations, as described in Section 4.2. The results reveal some notable behavioral differences compared to the in-sample setting. On the right-hand side of Figure 5, we report the aggregated rankings in the out-of-sample regime. Here, we can observe that Linear Regression and standard Decision Trees are adversely impacted by the out-of-sample setting, to the advantage of GLM, LASSO, and k-NN.
To better understand the origin of these differences, we examine the stratified results reported in Figure 8. The overall pattern remains consistent across most stratifications, with some deviations that complement the trends discussed in Section 6.1.1. For instance, k-NN tends to perform relatively better in the low-sample regime, while SR appears to be more adversely affected. When the number of features is small, Polynomial Regression, LASSO, and GLM show relative improvements. Conversely, k-NN is less negatively impacted by higher dimensionality, in contrast to its behavior under in-sample conditions. These tendencies are also reflected in metrics T2 and T3. Furthermore, GLM and LASSO tend to outperform Linear Regression as the linearity of the dataset decreases (metric L1). k-NN appears less affected by local noise (metric S3), while both GLM and LASSO achieve better performance when a feature exhibits strong correlation with the target variable (metric $\text{C1}_{\text{r}}$). In the same regime, Symbolic Regression faces a relative decay of its predictive performance.
Figure 8: Stratified results for the regression task (out-of-sample).
6.2.1 Training Time for Regression
Figure 9: Training times, measured in seconds, for the regression models. The horizontal dashed line represents the maximum allowed time for each method (300 seconds).
Figure 9 presents the training time of the evaluated models, a crucial aspect in scenarios requiring responsive model updates or rapid retraining.
Among the analyzed methods, IGANN exhibits the longest execution times, followed by EBM. Since IGANN relies on neural networks to compute shape functions, it would likely benefit from GPU acceleration. PR+LASSO shows the next highest training time, largely due to the expanded feature space created by the polynomial transformation. Interestingly, even GLM demonstrates relatively long training times despite its comparatively weak predictive performance.
Subsequently, k-NN, DT, LASSO, and Linear Regression show shorter training times. We emphasize the high variability in training times across methods, which highlights substantial differences in computational efficiency among interpretable models. A general trend also emerges: more expressive methods, those capable of representing richer classes of functions, tend to require significantly longer training times. Aside from GLM, methods such as EBM, IGANN, and PR+LASSO offer better predictive performance but demand one to two orders of magnitude more computation.
In Figure 9, Symbolic Regression is omitted since its execution time consistently reached the maximum threshold. Symbolic Regression requires extensive exploration of the functional space, a process that becomes particularly demanding for datasets with many samples, as each candidate expression must eventually be evaluated on the entire dataset (even with batching). This behavior also explains the underperformance of Symbolic Regression observed in the stratified analysis by sample size presented in Section 6.1.1.
7 Classification Results
7.1 Classification Performance (In-Sample)
Figure 10: Ranking comparison for the classification task between in-sample and out-of-sample settings.
For the classification task, we adopt the same analysis procedure used for regression, this time employing the F1 metric. Figure 10 reports the aggregated rankings computed across all 79 classification datasets; with the left-hand side displaying the in-sample results.
Across all classification datasets, three distinct groups of classifiers can be identified, although the separation between them is less pronounced than in the regression setting. EBM achieves the best performance by a clear margin, a trend that will also be observed in subsequent analyses. The second group comprises all remaining classifiers except Naive Bayes, which consistently performs the worst by a substantial margin.
It is noteworthy that traditional Decision Trees and GOSDT exhibit comparable predictive performance, with a slight advantage in favor of the former. Given that both methods produce similar model structures (i.e., decision trees), their interpretability can be directly compared by analyzing the size of the resulting trees. Figure 11 shows the distribution of the number of nodes in the retrieved models across all datasets. As expected, GOSDT yields significantly sparser trees, indicating more compact and interpretable models.
Figure 11: Comparison of model complexity, measured by tree size, between traditional Decision Trees and GOSDT.
7.1.1 Stratified Results for Classification
Figure 12: Stratified results for the classification task (in-sample).
As in Section 6.1.1, we stratify the classification results according to the dataset metrics introduced in Section 5.1. Figure 12 presents the stratified results for each metric.
Overall, the considered dataset metrics provide limited stratification of classifier performance (in contrast to the regression case), largely due to the smaller number of classification datasets (79 vs. 137). When stratifying into terciles or quartiles, the resulting strata contained too few datasets to yield meaningful comparisons. For this reason, we opted to stratify the classification datasets into halves, which provides a sample size per stratum comparable to that used in the regression analysis.
When considering the structural properties of the datasets (i.e., the number of rows, columns, and their ratio T2), we observe that the methods tend to perform similarly on datasets with a small number of samples or features. For datasets with a larger number of samples, Naive Bayes and GOSDT tend to perform worse, while k-NN and IGANN perform better, as expected. As the number of features increases, IGANN shows improved performance, which is expected given its neural network foundation, while decision-tree-based models, especially GOSDT, exhibit a decline in accuracy.
When examining the ratios T2 and T3, we observe the struggles of k-NN, already noticed in high dimensional feature space. This is consistent with its instance-based nature: as dimensionality grows, the feature space becomes increasingly sparse, and substantially more samples are needed to maintain good generalization. It is also noteworthy that Logistic Regression remains relatively robust both in low-sample settings and in high-dimensional scenarios.
For the dataset separability metric (F1), Logistic Regression and IGANN perform particularly well on highly separable datasets, with their performance declining as separability decreases. For the N3 metric (noisiness assessed via 1-NN error), classifiers tend to behave similarly when noise is high, suggesting noise reduces relative performance differences. Class entropy ($\text{C1}_{\text{c}}$) reveals that GOSDT is particularly sensitive to class imbalance. For the local class-entropy metric (E5), no strong overall trend is visible beyond the weaker performance of decision trees at lower E5 values.
7.2 Out-Of-Sample Classification
On the right-hand side of Figure 10, we report the aggregated results for the out-of-sample regime. It is worth noting that the out-of-sample setting tends to increase performance variability thus reducing the relative differences across methods.
The stratified results, shown in Figure 13, display trends consistent with those observed in the in-sample analysis. Two remarkable deviations emerge from this comparison. First, when considering the number of features, Logistic Regression exhibits a relative improvement as the dimensionality increases. Second, the sensitivity of k-NN to structural metrics is significantly exacerbated in the out-of-sample setting. Specifically, k-NN shows markedly lower performance on low-sample datasets, high-dimensional datasets, and datasets characterized by an unfavorable ratio of rows to columns. This pattern strongly suggests that k-NN is particularly vulnerable when there are insufficient out-of-sample examples to adequately describe or tile the feature space, leading to less reliable neighborhood identification. Furthermore, when assessing the feature separability (metric F1), k-NN tends to underperform for lower F1 values, that is, when a specific feature is highly effective at discriminating between classes, indicating that its reliance on local proximity cannot represent simple, global linear separators that other models, like Logistic Regression, exploit effectively.
Figure 13: Stratified results for the classification task (out-of-sample).
7.2.1 Training Time for Classification
Figure 14: Training times, measured in seconds, for the classification models. The horizontal dashed line represents the maximum allowed time for each method (300 seconds).
The training time evaluation for the classification task follows the same procedure as described for regression. Figure 14 presents the training times of the classification models.
As observed in the regression task, most of the observed results align with expectations. IGANN exhibits the longest training time, often approaching the maximum allowed duration of 300 seconds during final training. GOSDT shows the second-highest training time, approximately two orders of magnitude greater than that of ordinary Decision Trees, which are among the fastest methods. EBM and, unexpectedly, Logistic Regression follow, while k-NN, Decision Trees, and Naive Bayes achieve the shortest training times overall.
Notably, Logistic Regression, despite sharing the same linear formulation, trains slower than Linear Regression in the regression task. This difference arises because Logistic Regression relies on gradient descent optimization, whereas Linear Regression uses the closed-form Ordinary Least Squares solution.
Moreover, an interesting trade-off emerges between interpretability and training time: more interpretable or sparse models, such as IGANN and GOSDT, tend to require longer training times compared to methods like EBM and Decision Trees. This additional computational cost arises from the need to search for sparser solutions, as in GOSDT, or to optimize smoother shape functions, as in IGANN.
8 Discussion
Our results can be contextualized by examining previous findings in the literature. The work of kraus2023interpretable, who introduced IGANN, also utilized a stratified analysis on PMLB datasets (categorized by sample size and dimensionality). They showed that IGANN could outperform other interpretable models, including EBM, especially on small or low-dimensional data. This contrasts with our findings that EBM often exhibits stronger predictive performance, a difference that we can attribute to the limited hyperparameter search space used in that study for EBM (where the number of possible interactions was set to zero). While this can favor interpretability, it also potentially affects predictive performance. We also observe that, coherently with that study, the execution time of IGANN is significantly worse compared to EBM and other competing models. Our classification results align with theirs, showing EBM generally outperforms IGANN in that task.
kruschel2025challenging evaluated a set of interpretable models on 20 datasets, noting that while black-box models performed the best in 14 out of 20 cases, the performance gap with strong interpretable models (especially from the GAM family) was small. They also attempted to measure model interpretability by using some criteria introduced by sudjianto2021designing, which yielded counter-intuitive results. For instance, they place Decision Trees on a low position on their interpretability scale, despite their widespread acceptance as being highly interpretable. This ambiguity around objective interpretability metrics supports our decision to avoid comparing interpretability, with the only exceptions being decision-tree-based and linear models, since they share the same model structure.
Consistent with the findings of la2021contemporary and wilstrup2021symbolic, our evaluation confirms that Symbolic Regression consistently outperforms traditional Linear Regression models. wilstrup2021symbolic, who tested SR on 48 PMLB datasets, specifically concluded that SR models generalize better to out-of-sample data than Linear Regression, LASSO, and Decision Trees. This evidence aligns with our observation that SR captures non-linear relationships that simple linear models miss, making it a stronger alternative to LR.
Comparisons involving traditional classification families also appear in wainer2016comparison, who analyzed 14 algorithm families on 115 datasets. They found that k-NN generally outperforms Naive Bayes, with GLMNET (related to our Logistic Regression/LASSO models) positioning between the two. This provides a reference point for the ranking of the simpler models in our study, which is consistent with their findings.
Finally, the work of scholz_comparison_2021 provides a relevant contrast due to their stratified analysis on synthetic data based on the structural properties of the data (number of rows and columns) and relations with the target variable. They found that no classifier always dominates and that performance varies across scenarios. Unlike our findings, they concluded that Logistic Regression is fast but not competitive in predictive accuracy; we observed that Logistic Regression remains a competitive option in specific scenarios, such as those involving high dimensionality, scarce data, or the presence of a highly discriminative feature. Consistent with our results, they found that LASSO-type regularization often does not improve performance enough to be a top choice and that individual Decision Tree models are generally not among the strongest performers. A notable divergence, however, concerns k-NN. scholz_comparison_2021 found that k-NN surprisingly achieved a high rank in small, high-dimensional synthetic datasets. In contrast, our analysis on real-world data highlights this specific regime as a weakness for k-NN, particularly in out-of-sample settings where the feature space is not sufficiently tiled by training examples.
9 Conclusions
This study presented an extensive empirical evaluation of intrinsically interpretable machine learning models, addressing the lack of systematic guidance for practitioners in the field. By benchmarking sixteen diverse methods across more than two hundred real-world datasets, we moved beyond aggregate performance metrics to uncover how model behavior is influenced by structural data characteristics such as sample size, dimensionality, linearity, and class imbalance. Our analysis confirms that while modern interpretable approaches can achieve competitive predictive accuracy, model selection should be carefully tailored to the specific properties of the data at hand rather than relying on a one-size-fits-all solution.
In the regression domain, we observed a clear performance hierarchy where EBM consistently demonstrated superior predictive accuracy. However, the stratification analysis revealed distinct niches for other methods. Symbolic Regression and IGANN proved robust in non-linear regimes, whereas traditional linear and regularized models remained effective choices for high-dimensional or linear problems. The distinction was less pronounced in classification tasks, where performance tiers were narrower. Nevertheless, EBM retained a general advantage, while GOSDT offered a compelling trade-off, delivering significantly sparser and more interpretable models than standard Decision Trees at the cost of increased training time. Our investigation into out-of-sample generalization further highlighted critical vulnerabilities in specific methods, such as the degradation of k-NN in high-dimensional spaces under distributional shifts, contrasting with the robustness of Logistic Regression in data-scarce or high-dimensional scenarios.
A significant insight emerging from this study is the tangible cost associated with interpretability. We found that models offering higher expressiveness or provable sparsity, such as IGANN, Symbolic Regression, and GOSDT, require training times that are orders of magnitude higher than their simpler counterparts. This computational overhead presents a practical barrier for real-time applications or large-scale data streams.
In this work, we deliberately focused on predictive performance and computational efficiency, and did not attempt to quantify interpretability itself, reflecting the absence of a universally accepted and objective interpretability metric. Developing rigorous, model-agnostic frameworks for measuring interpretability, therefore, remains an important direction for future research.
Appendix A Classification and Regression Methods
In this study, we evaluate a diverse set of interpretable machine learning models for both classification and regression tasks. These methods are selected to represent a range of modeling paradigms, from linear and tree-based approaches to probabilistic, instance-based, and symbolic methods. We focus on models that are intrinsically interpretable and whose implementations are readily available and stable. The following subsections provide an overview of each model category, highlighting their interpretability characteristics and core assumptions.
A.1 Linear models: Logistic, Linear, Lasso, GLM, GAM
Linear models are among the most widely used interpretable methods due to their simplicity and transparency. Linear Regression models continuous outcomes as a linear combination of input features. In classification tasks, Logistic Regression estimates the probability of categorical outcomes by applying a logistic function to a linear combination of features, mapping logit values to probabilities.
Extensions of linear models, such as LASSO regression, introduce regularization to promote sparsity, improving interpretability by selecting a subset of relevant features through an L1 penalty (tibshirani1996regression). Generalized Linear Models extend linear models to accommodate non-normal target distributions (e.g., Poisson, Gamma) via appropriate link functions (mccullagh2019generalized). Generalized Additive Models (GAMs) further enhance flexibility by allowing non-linear transformations of individual features, such as splines, or interactions between feature pairs, while maintaining an additive structure that preserves interpretability (hastie1986generalized). Different formulations and implementations of GAMs have been proposed. Explainable Boosting Machines (EBMs) extend the concept of GAMs using gradient boosting to capture complex feature interactions while preserving interpretability through per-feature visualizations and monotonic constraints (nori2019interpretml).
Recently, kraus2023interpretable introduced a novel variant, the Interpretable Generalized Additive Neural Network (IGANN), which promotes linearity by initializing shape functions linearly and allowing deviations only when required by the underlying features.
A.2 Tree-based models: Decision Tree, GOSDT
Tree-based models provide a rule-based representation of decision processes. Decision Trees recursively partition the feature space into homogeneous regions, producing hierarchical rules that are easy to visualize (breiman1984classification). However, their interpretability and predictive performance can be compromised by suboptimal or overly large trees (lipton2018mythos), as classical algorithms often rely on greedy splitting heuristics. Optimal Sparse Decision Trees (OSDT) introduced a practical, bound-driven search for provably optimal trees on binary variables, improving sparsity compared to greedy methods (hu2019optimal). Subsequent work generalized and scaled this approach to handle continuous features and imbalanced data (lin2020generalized). More recent methods apply reference ensembles from black-box models to guide the search, substantially reducing computation while producing sparse trees that remain interpretable (mctavish2022fast).
A.3 Probabilistic models: Naive Bayes
Naive Bayes is a probabilistic approach grounded in Bayes’ theorem, assuming conditional independence among features (hastie2009elements). Despite its simplicity, it is effective in many real-world scenarios and provides interpretable probabilistic outputs. Feature contributions to predictions can be directly assessed through the model’s likelihood functions, offering insights into how individual variables influence the predicted probabilities.
A.4 Instance-based: KNN
Instance-based methods, such as k-Nearest Neighbors (k-NN), make predictions by comparing new instances to stored examples in the training data (cover1967knn). k-NN is often considered interpretable because predictions can be traced back to the most similar observed instances, providing an intuitive understanding of how the model makes decisions.
While some may argue that instance-based methods offer only local explanations, similar to post-hoc methods like LIME (ribeiro2016should), k-NN’s predictions are fully determined by the nearest training samples, without any additional approximation or model transformation. Consequently, the behavior of the model across the entire dataset can, in principle, be understood by examining the stored instances, and the influence of each feature is directly observable through the similarity metric.
A.5 Symbolic models: Symbolic Regression
Symbolic Regression seeks to discover explicit mathematical expressions that describe the relationship between input features and the target variable (la2021contemporary). By generating human-readable formulas, Symbolic Regression offers a high degree of interpretability, allowing domain experts to understand and validate the underlying patterns captured by the model. Methods in this category balance predictive accuracy with simplicity, often guided by constraints on expression complexity.
Appendix B Hyperparameter Ranges
Table 3 reports the hyperparameter ranges used for model selection across all interpretable methods. Search spaces were chosen to reflect standard practice while enforcing constraints that favor parsimonious and interpretable model configurations.
Model Tuning Parameters Range
k-NN (clf/regr) n_neighborsn\_neighbors $3,\min(100,0.2\cdot n\_{samples})]$
Decision Tree (clf) min_samples_leafmin\_samples\_leaf $[0.01,0.3]$
criterioncriterion {gini, entropy}
max_depthmax\_depth $2,\min(6,n\_{features}/2+1)]$
Decision Tree (regr) max_depthmax\_depth $2,\min(6,n\_{features}/2+1)]$
Logistic Regression $C$ $10^{-2},10^{1}]$ (log-scale)
penaltypenalty {l2, None}
GOSDT GDBT_n_estimatorsGDBT\_n\_estimators $[20,40]$
GDBT_max_depthGDBT\_max\_depth $[2,5]$
GOSDT_regularizationGOSDT\_regularization $[0.01,0.2]$ (log-scale)
GOSDT_similar_supportGOSDT\_similar\_support {False, True}
GOSDT_depth_budgetGOSDT\_depth\_budget $[2,6]$
Naive Bayes var_smoothingvar\_smoothing $10^{-12},10^{-1}]$ (log-scale)
EBM (clf/regr) interactionsinteractions $[0,2]$
learning_ratelearning\_rate $[0.01,0.05]$ (log-scale)
max_roundsmax\_rounds $[100,200]$
min_samples_leafmin\_samples\_leaf $[2,5]$
IGANN (clf/regr) init_reginit\_reg $10^{-7},1]$ (log-scale)
n_hidn\_hid $[5,20]$
elm_alphaelm\_alpha $10^{-7},1]$ (log-scale)
elm_scaleelm\_scale $[1,10]$
boost_rateboost\_rate $[0.1,1]$
n_estimatorsn\_estimators $[100,5000]$
Symbolic Regression select_k_featuresselect\_k\_features $2,\min(n\_{features},12)]$
operatorsoperators $\{+,-,\times,/,exp,\log,\sin,\cos\}$
populationspopulations $[6,20]$
population_sizepopulation\_size $[32,96]$
weight_optimizeweight\_optimize $10^{-3},10^{-1}]$ (log-scale)
parsimonyparsimony $10^{-3},1]$ (log-scale)
maxsizemaxsize $[15,70]$
batchingbatching AutoAuto
Linear Regression None None
Lasso alphaalpha $10^{-4},10^{2}]$ (log-scale)
GLM alphaalpha $10^{-4},10^{1}]$ (log-scale)
Polynomial Regr. + Lasso poly__degreepoly\_\_degree {2, 3}
alphaalpha $10^{-4},10^{2}]$ (log-scale)
Table 3: Hyperparameter search spaces for classification and regression models.